OpenStudy (anonymous):

Given the equation y − 4 =3/4 (x + 8) in point-slope form, identify the equation of the same line in slope-intercept form.

hartnn (hartnn):

isolate y by adding 4 on both sides... because the slope-point form is y=mx+c

OpenStudy (anonymous):

so would the answer be y=3/4x+10

hartnn (hartnn):

very correct :) good.....

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

i need help on another one though

hartnn (hartnn):

sure,go on....

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

hartnn (hartnn):

first find the slope using m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

well i got -1/-1

hartnn (hartnn):

whats x2-x1??

OpenStudy (anonymous):

-3-2=-1

hartnn (hartnn):

x1 is -3,x2 is 2 so x2-x1=2-(-3)=2+3=5......got this??

OpenStudy (anonymous):

i got it already

hartnn (hartnn):

u got final answer also?